# Complete non-ambiguous trees and associated permutations: new enumerative results

### v2

@@ 4.2. Linking permutations with two and three CNATs
Since $\pi$ contains the pattern $\tau$, the permutation graph $G_{\pi}$ contains an induced subgraph graph $H$ which is isomorphic to $G_{\tau}$ by the remarks at the end of Section 2.3. Note that we have $t_{G_{\tau}}=t_{H}$ in this case. The graph $G_{\pi}$ can then be constructed as follows. Fix some spanning tree $G^{\prime}$ of $G_{\pi}$. Let $G^{0}$ be the graph with vertex set $[n]=V\left(G_{\pi}\right)$ and edge set $E\left(G^{0}\right):=E(G^{\prime})\cup E(H)$. By construction, since $G^{\prime}$ is a tree, we have $\mathrm{Core}_{2}\left([\right)G^{0}]=\mathrm{Core}_{2}\left([\right)H]$, so that $t_{G^{0}}=t_{H}$ by Lemma 2.5. Moreover, we have $E\left(G^{0}\right)\subseteq E\left(G_{\pi}\right)$, so that the graph $G_{\pi}$ can be obtained from $G^{0}$ through a (possibly empty) sequence of edge additions. Applying Lemma 2.7 then yields that $t_{G^{0}}\leq t_{G_{\pi}}$. Combined with the above, we get $t_{G_{\tau}}=t_{H}\leq t_{G_{\pi}}$, and the result follows from Theorem 2.12. ∎
@@
First suppose that $\pi$ contains a unique occurrence of the $321$ pattern, and avoids $3412$. By Proposition 2.10, this implies that the permutation graph $G_{\pi}$ of $G$ is a decorated $3$-cycle. From Lemma 2.6, we then obtain $t_{G}=3-1=2$, which implies by Theorem 2.12, that $\pi$ has exactly $2$ CNATs.
Now suppose that $\pi$ has exactly $2$ CNATs. We know from Lemma 4.9 that $\pi$ has a (unique) fixed point $j\in\{2,\cdots,n-1\}$. Consider the lower-left $j$-quadrant $\Pi_{<j,\geq j}$. By Proposition 4.10 it has a unique dot, and this dot cannot be in row $j$, since the dot in row $j$ is in column $j$. Thus this dot must be $(i,\pi_{i})$ with $i<j$ and $\pi_{i}>j=\pi_{j}$. Similarly, the upper-right $j$-quadrant $\Pi_{\geq j,<j}$ has a unique dot which is not in column $j$, so is therefore $(k,\pi_{k})$ with $k>j$ and $\pi_{k}<j=\pi_{j}$. In other words, $\pi_{i},\pi_{j}=j,\pi_{k}$ forms a $321$ pattern in the permutation $\pi$, or equivalently induces a $3$-cycle in the permutation graph $G_{\pi}$.
We show that there can be no other occurrences of $321$, or any occurrence of $3412$, in the permutation $\pi$. Otherwise, $G_{\pi}$ would contain another cycle other than that induced on $\pi_{i},\pi_{j},\pi_{k}$. In particular this would imply that the $2$-core $\mathrm{Core}_{2}\left([\right)G_{\pi}]$ is not isomorphic to the $3$-cycle, and therefore we would have $t_{G_{\pi}}>t_{C_{3}}=2$ by Lemma 2.8. This completes the proof of the proposition. ∎
The $321$ pattern structure for $\pi\in B(n,2)$ can be conveniently represented using the terminology of mesh patterns. Mesh patterns were introduced by Brändén and Claesson [9] as a generalisation of classical permutation patterns with added restrictions. Formally a mesh pattern is a pair $p=(\tau,R)$, with $\pi\in S_{k}$ and $R\subseteq[0,k]\times[0,k]$ for some positive integer $k$, where $[0,k]$ denotes the integer interval $\{0,\cdots,k\}$. To depict a mesh pattern, we plot the permutation $\tau$ on a $k\times k$ grid (with our standard directions from top to bottom on the rows, and left to right on the columns), and for each $(i,j)\in R$, we shade the unit with top-left corner $(i,j)$. For example, the mesh pattern $p=\left(3241,\{(0,2),(1,3),(1,4),(4,2),(4,3)\}\right)$ is depicted on Figure 11 below.
11112222333344441111222233334444
@@
We then say that a permutations $\pi$ contains a mesh pattern $p=(\tau,R)$ if $\pi$ contains the (classical) pattern $\tau$ in such a way that the shaded regions do not contain any elements of $\pi$. For a more formal definition, see [9]. For example, the permutation $\pi=53241$ contains the mesh pattern $p$ from Figure 11, as the $5$ is not in a forbidden region. However, the permutation $\pi=35241$ does not contain $p$, as the $5$ is in the forbidden region between columns $1$ and $2$ on Figure 11 (and there is only one occurrence of $\tau=3241$ in $\pi$). With this terminology, a permutation $\pi\in B(n,2)$ contains a (unique) occurrence of the mesh pattern depicted in Figure 12 below. To check this, note that if any of the shaded regions contained a dot, we would get a second occurrence of the (classical) $321$ pattern. For example, if any region to the left of column $j$ and below row $j$ contained a dot, this would form an occurrence of $321$ with $\pi_{j}$ and $\pi_{k}$.
$i$$j$$k$$\pi_{k}$$\pi_{j}=j$$\pi_{i}$
@@
First suppose that $\pi$ contains a unique occurrence of the $3412$ pattern, and avoids $321$. By Proposition 2.10, this implies that the permutation graph $G_{\pi}$ of $G$ is a decorated $4$-cycle. From Lemma 2.6, we then obtain $t_{G}=4-1=3$, which implies by Theorem 2.12, that $\pi$ has exactly $3$ CNATs.
@@
First, note that if $G$ is a tree, then $\mathrm{cnat}\left([\right)\pi]=t_{G}=1\neq 3$ (by Lemma 2.4), so that $G$ must contain at least one cycle. If $G$ is a decorated $3$-cycle, then by Lemma 2.6 we would have $\mathrm{cnat}\left([\right)\pi]=t_{G}=2\neq 3$. Therefore $G$ either contains at least two cycles, or a cycle of length at least $4$. If $G$ contains a cycle of length $k\geq 5$, then Lemma 2.8 implies that $\mathrm{cnat}\left([\right)\pi]=t_{G}\geq t_{C_{k}}=k-1\geq 4$, a contradiction. Similarly, if $G$ contains a $4$-cycle, to get $t_{G}=3$ the $2$-core $\mathrm{Core}_{2}\left([\right)G]$ must be isomorphic to $C_{4}$ (again using Lemma 2.8), i.e. $G$ is a decorated $4$-cycle.
It therefore remains to show that $G$ cannot contain two $3$-cycles. We again seek contradiction. If $G$ has two $3$-cycles which share an edge, then the “outer cycle” is a $4$-cycle (see Figure 17(a)). One can check that this graph $G^{\prime}$ satisfies $t_{G^{\prime}}=4$, and again by Lemma 2.8 we get the contradictory $\mathrm{cnat}\left([\right)\pi]=t_{G}\geq t_{G^{\prime}}=4$. It therefore remains to consider the case where $G$ contains at least two edge-disjoint $3$-cycles. Since $G$ is connected, these are necessarily joined in $G$ by a path of $k$ edges for some $k\geq 0$. Let $G^{\prime}$ be the subgraph of $G$ consisting of the two $3$-cycles together with this joining path (see Figure 13). It is straightforward to check that we have $t_{G^{\prime}}=4$, and the result follows as above. This completes the proof of our claim that if $\pi$ has exactly $3$ CNATs, the associated permutation graph $G_{\pi}$ is a decorated $4$-cycle, as desired.
@@
We can directly construct the three CNATs corresponding to the permutation $\pi$ based on the $3412$ pattern of Figure 14. The construction is similar to that of Figure 9. Namely, we define three cells $c:=(j,\pi_{\ell}),c_{\mathrm{left}}:=(i,\pi_{\ell}),c_{\mathrm{up}}:=(j,\pi_{k})$. Deleting the dots $(j,\pi_{j})$ and $(\ell,\pi_{\ell})$ from the permutation $\pi$ yields a permutation $\pi^{\prime}$ which has a single CNAT $T^{\prime}$ (the permutation graph $G_{\pi^{\prime}}$ is acyclic, since $\pi^{\prime}$ avoids $321$ and $3412$) [^3] As in Remark 4.8 we are slightly abusive here, since we need to take care not to disconnect the permutation graph.. The three CNATs associated with $T$ are then constructed from $T^{\prime}$ by adding dots in any two of the three cells $c=(j,\pi_{\ell}),\ c_{\mathrm{left}}=(i,\pi_{\ell}),\ c_{\mathrm{up}}=(j,\pi_{k})$, as on Figure 9.
@@
    There exists a unique value of $k\in\{2,\cdots n\}$ such that the lower-left and upper-right $k$-quadrants of $\Pi$ have exactly two dots.
@@
With Propositions 4.12 and 4.13, we are now equipped to define the bijection between $B(n,2)$ and $B(n+1,3)$. Let $\pi\in B(n,2)$ be an $n$-permutation with $2$ CNATs, and suppose that $\pi_{i},\pi_{j}=j,\pi_{k}$ is the unique occurrence of the pattern $321$ in $\pi$, with $i<j<k$. We first define an $(n+1)$-permutation $\tilde{\pi}$ by $\tilde{\pi}:=\mathrm{Insert}_{[}j+1](\pi)$. Since $\pi$ has a unique occurrence of the pattern $321$, given by $\pi_{i},\pi_{j}=j,\pi_{k}$, it follows by construction that $\tilde{\pi}$ has a unique occurrence of the pattern $4231$, given by $\tilde{\pi}_{i}=\pi_{i}+1,\tilde{\pi}_{j}=j,\tilde{\pi}_{j+1}=j+1,\tilde{\pi}_{k+1}=\pi_{k}$. Moreover, by construction, $\tilde{\pi}$ has exactly two occurrences of $321$, given by $\tilde{\pi}_{i},\tilde{\pi}_{j},\tilde{\pi}_{k+1}$ and $\tilde{\pi}_{i},\tilde{\pi}_{j+1},\tilde{\pi}_{k+1}$. We also claim that $\tilde{\pi}$ avoids $3412$. Indeed, since $\pi$ avoids $3412$, the only way the $3412$ pattern could occur in $\tilde{\pi}$ is by using both $\tilde{\pi}_{j}=j$ and $\tilde{\pi}_{j+1}=j+1$. This means that $\tilde{\pi}_{j}\tilde{\pi}_{j+1}$ must be either the $34$ or the $12$ part of the $3412$ pattern. But if it were the $34$ part, then $\pi_{i}\pi_{j}x$ could be completed into two occurrences of $321$ in $\pi$ by setting $x$ to be the $1$ or the $2$ in the $3412$ occurrence in $\tilde{\pi}$ (these are the same in $\pi$ and $\tilde{\pi}$). Similarly, if $\tilde{\pi}_{j}\tilde{\pi}_{j+1}$ were the $12$ part of the $3412$ occurrence in $\tilde{\pi}$, then $y\pi_{j}\pi_{k}$ could be completed into two occurrences of $321$ in $\pi$ by setting $y$ to be the $3$ or the $4$ in the $3412$ occurrence in $\tilde{\pi}$. In both cases this contradicts the fact that $\pi$ contains a unique occurrence of $321$, and so the claim that $\tilde{\pi}$ avoids $3412$ is proved. We then define a permutation $\pi^{\prime}$, obtained from $\tilde{\pi}$ by changing the $4231$ pattern to a $3412$ pattern (on the same elements), and leaving all other elements unchanged. More formally, we have:
(5) $\begin{cases}\pi^{\prime}_{i}=\tilde{\pi}_{j+1}=j+1\\ \pi^{\prime}_{j}=\tilde{\pi}_{i}=\pi_{i}+1\\ \pi^{\prime}_{j+1}=\tilde{\pi}_{k+1}=\pi_{k}\\ \pi^{\prime}_{k+1}=\tilde{\pi}_{j}=j\\ \pi^{\prime}_{\ell}=\tilde{\pi}_{\ell}\qquad\text{for all other values of }\ell\end{cases}.$
@@
The construction $\pi\mapsto\pi^{\prime}$ is illustrated on Figure 15 below. On the left (Figure 14(a)) we have the occurrence $\pi_{i},\pi_{j}=j,\pi_{k}$ of $321$ in the permutation $\pi\in B(n,2)$. On the right (Figure 14(b)) we have the occurrence $\pi^{\prime}_{i}=\pi_{j}+1=j+1,\pi^{\prime}_{j}=\pi_{i}+1,\pi^{\prime}_{j+1}=\pi_{k},\pi^{\prime}_{k+1}=\pi_{j}=j$ of $3412$ in $\pi^{\prime}$. The other dots of $\pi$ are unchanged in $\pi^{\prime}$, after the necessary shifting downwards and/or to the right to create a permutation.
$i$$j$$k$$\pi_{k}$$\pi_{j}=j$$\pi_{i}$
@@
Let $\pi\in B(n,2)$. By construction $\pi^{\prime}$ is an $(n+1)$-permutation. We first show that $\pi^{\prime}\in B(n+1,3)$, using Proposition 4.13. We first refine the construction of Figure 15, using the mesh pattern of $\pi$ illustrated in Figure 12. For convenience, we represent this as a block decompositions of the permutations $\pi$ and $\pi^{\prime}$, as illustrated on Figure 16 below. The unlabelled blocks correspond to the shaded regions of the mesh patterns, and are therefore empty.

### v4

@@ 4.2. Linking permutations with two and three CNATs
Since $\pi$ contains the pattern $\tau$, the permutation graph $G_{\pi}$ contains an induced subgraph $H$ which is isomorphic to $G_{\tau}$ by the remarks at the end of Section 2.3. Note that we have $a_{G_{\tau}}=t_{H}$ in this case. The graph $G_{\pi}$ can then be constructed as follows. Fix some spanning tree $G^{\prime}$ of $G_{\pi}$. Let $G^{0}$ be the graph with vertex set $n]=V\left(G_{\pi}\right)$ and edge set $E\left(G^{0}\right):=E(G^{\prime})\cup E(H)$. By construction, since $G^{\prime}$ is a tree, we have $\mathrm{Core}_{2}\left([\right)G^{0}]=\mathrm{Core}_{2}\left([\right)H]$, so that $a_{G^{0}}=t_{H}$ by Lemma 2.5. Moreover, we have $E\left(G^{0}\right)\subseteq E\left(G_{\pi}\right)$, so that the graph $G_{\pi}$ can be obtained from $G^{0}$ through a (possibly empty) sequence of edge additions. Applying Lemma 2.7 then yields that $a_{G^{0}}\leq a_{G_{\pi}}$. Combined with the above, we get $a_{G_{\tau}}=t_{H}\leq a_{G_{\pi}}$, and the result follows from Theorem 2.13. ∎
@@
First suppose that $\pi$ contains a unique occurrence of the $321$ pattern, and avoids $3412$. By Proposition 2.10, this implies that the permutation graph $G_{\pi}$ of $\pi$ is a decorated $3$-cycle. From Lemma 2.6, we then obtain $a_{G}=3-1=2$, which implies by Theorem 2.13, that $\pi$ has exactly $2$ CNATs.
Now suppose that $\pi$ has exactly $2$ CNATs. We know from Lemma 4.8 that $\pi$ has a (unique) fixed point $j\in\{2,\ldots,n-1\}$. Consider the lower-left $j$-quadrant $\Pi_{<j,\geq j}$. By Proposition 4.9 it has a unique dot, and this dot cannot be in row $j$, since the dot in row $j$ is in column $j$. Thus this dot must be $(i,\pi_{i})$ with $i<j$ and $\pi_{i}>j=\pi_{j}$. Similarly, the upper-right $j$-quadrant $\Pi_{\geq j,<j}$ has a unique dot which is not in column $j$, so is therefore $(k,\pi_{k})$ with $k>j$ and $\pi_{k}<j=\pi_{j}$. In other words, $\pi_{i},\pi_{j}=j,\pi_{k}$ forms a $321$ pattern in the permutation $\pi$, or equivalently induces a $3$-cycle in the permutation graph $G_{\pi}$.
We show that there can be no other occurrences of $321$, or any occurrence of $3412$, in the permutation $\pi$. Otherwise, $G_{\pi}$ would contain another cycle other than that induced on $\pi_{i},\pi_{j},\pi_{k}$. In particular this would imply that the $2$-core $\mathrm{Core}_{2}\left([\right)G_{\pi}]$ is not isomorphic to the $3$-cycle, and therefore we would have $a_{G_{\pi}}>a_{C_{3}}=2$ by Lemma 2.8. This completes the proof of the proposition. ∎
The $321$ pattern structure for $\pi\in B(n,2)$ can be conveniently represented using the terminology of mesh patterns. Mesh patterns were introduced by Brändén and Claesson [9] as a generalisation of classical permutation patterns with added restrictions. Formally a mesh pattern is a pair $p=(\tau,R)$, with $\tau\in S_{k}$ and $R\subseteq[0,k]\times[0,k]$ for some positive integer $k$, where $[0,k]$ denotes the integer interval $\{0,\ldots,k\}$. To depict a mesh pattern, we plot the permutation $\tau$ on a $k\times k$ grid (with our standard directions from top to bottom on the rows, and left to right on the columns), and for each $(i,j)\in R$, we shade the unit with top-left corner $(i,j)$. For example, the mesh pattern $p=\left(3241,\{(0,2),(1,3),(1,4),(4,2),(4,3)\}\right)$ is depicted in Figure 11 below.
$11223344$
@@
We then say that a permutation $\pi$ contains a mesh pattern $p=(\tau,R)$ if $\pi$ contains the (classical) pattern $\tau$ in such a way that the shaded regions do not contain any elements of $\pi$. For a more formal definition, see [9]. For example, the permutation $\pi=53241$ contains the mesh pattern $p$ from Figure 11, as the $5$ is not in a forbidden region. However, the permutation $\pi=35241$ does not contain $p$, as the $5$ is in the forbidden region between columns $1$ and $2$ in Figure 11 (and there is only one occurrence of $\tau=3241$ in $\pi$). With this terminology, a permutation $\pi\in B(n,2)$ contains a (unique) occurrence of the mesh pattern depicted in Figure 12 below. To check this, note that if any of the shaded regions contained a dot, we would get a second occurrence of the (classical) $321$ pattern. For example, if any region to the left of column $j$ and below row $j$ contained a dot, this would form an occurrence of $321$ with $\pi_{j}$ and $\pi_{k}$.
[…]
@@
First suppose that $\pi$ contains a unique occurrence of the $3412$ pattern, and avoids $321$. By Proposition 2.10, this implies that the permutation graph $G_{\pi}$ of $\pi$ is a decorated $4$-cycle. From Lemma 2.6, we then obtain $a_{G_{\pi}}=4-1=3$, which implies by Theorem 2.13, that $\pi$ has exactly $3$ CNATs.
@@
First, note that if $G$ is a tree, then $\mathrm{cnat}\left([\right)\pi]=a_{G}=1\neq 3$ (by Lemma 2.4), so that $G$ must contain at least one cycle. If $G$ is a decorated $3$-cycle, then by Lemma 2.6 we would have $\mathrm{cnat}\left([\right)\pi]=a_{G}=2\neq 3$. Therefore $G$ either contains at least two cycles, or a cycle of length at least $4$. If $G$ contains a cycle of length $k\geq 5$, then Lemma 2.8 implies that $\mathrm{cnat}\left([\right)\pi]=a_{G}\geq a_{C_{k}}=k-1\geq 4$, a contradiction. Similarly, if $G$ contains a $4$-cycle, to get $a_{G}=3$ the $2$-core $\mathrm{Core}_{2}\left([\right)G]$ must be isomorphic to $C_{4}$ (again using Lemma 2.8), i.e. $G$ is a decorated $4$-cycle.
It therefore remains to show that $G$ cannot contain two $3$-cycles. We again seek contradiction. If $G$ has two $3$-cycles which share an edge, then the “outer cycle” is a $4$-cycle (see Figure 18(a)). One can check that this graph $G^{\prime}$ satisfies $a_{G^{\prime}}=4$, and again by Lemma 2.8 we get the contradictory $\mathrm{cnat}\left([\right)\pi]=a_{G}\geq a_{G^{\prime}}=4$. It therefore remains to consider the case where $G$ contains at least two edge-disjoint $3$-cycles. Since $G$ is connected, these are necessarily joined in $G$ by a path of $k$ edges for some $k\geq 0$. Let $G^{\prime}$ be the subgraph of $G$ consisting of the two $3$-cycles together with this joining path (see Figure 13). It is straightforward to check that we have $a_{G^{\prime}}=4$, and the result follows as above. This completes the proof of our claim that if $\pi$ has exactly $3$ CNATs, the associated permutation graph $G_{\pi}$ is a decorated $4$-cycle, as desired.
@@
We can directly construct the three CNATs corresponding to the permutation $\pi$ based on the $3412$ pattern of Figure 14. The construction is similar to that of Figure 9. Namely, we define three cells $c:=(j,\pi_{\ell}),c_{\mathrm{left}}:=(i,\pi_{\ell}),c_{\mathrm{up}}:=(j,\pi_{k})$. Deleting the dots $(j,\pi_{j})$ and $(\ell,\pi_{\ell})$ from the permutation $\pi$ yields a permutation $\pi^{\prime}$ which has a single CNAT $T^{\prime}$ (the permutation graph $G_{\pi^{\prime}}$ is acyclic, since $\pi^{\prime}$ avoids $321$ and 34123412) [^3] As in Proof of Lemma 4.7 we are slightly abusive here, since we need to take care not to disconnect the permutation graph.. The three CNATs associated with $T$ are then constructed from $T^{\prime}$ by adding dots in any two of the three cells $c=(j,\pi_{\ell}),\ c_{\mathrm{left}}=(i,\pi_{\ell}),\ c_{\mathrm{up}}=(j,\pi_{k})$, as in Figure 9.
@@
    There exists a unique value of $k\in\{2,\ldots,n\}$ such that the lower-left and upper-right $k$-quadrants of $\Pi$ have exactly two dots.
@@
With Propositions 4.11 and 4.12, we are now equipped to define the bijection between $B(n,2)$ and $B(n+1,3)$. Let $\pi\in B(n,2)$ be an $n$-permutation with $2$ CNATs, and suppose that $\pi_{i},\pi_{j}=j,\pi_{k}$ is the unique occurrence of the pattern $321$ in $\pi$, with $i<j<k$. We first define an $(n+1)$-permutation $\tilde{\pi}$ by $\tilde{\pi}:=\mathrm{Insert}_{[}j+1](\pi)$. Since $\pi$ has a unique occurrence of the pattern $321$, given by $\pi_{i},\pi_{j}=j,\pi_{k}$, it follows by construction that $\tilde{\pi}$ has a unique occurrence of the pattern $4231$, given by $\tilde{\pi}_{i}=\pi_{i}+1,\tilde{\pi}_{j}=j,\tilde{\pi}_{j+1}=j+1,\tilde{\pi}_{k+1}=\pi_{k}$. Moreover, by construction, $\tilde{\pi}$ has exactly two occurrences of $321$, given by $\tilde{\pi}_{i},\tilde{\pi}_{j},\tilde{\pi}_{k+1}$ and $\tilde{\pi}_{i},\tilde{\pi}_{j+1},\tilde{\pi}_{k+1}$. We also claim that $\tilde{\pi}$ avoids $3412$. Indeed, since $\pi$ avoids $3412$, the only way the $3412$ pattern could occur in $\tilde{\pi}$ is by using both $\tilde{\pi}_{j}=j$ and $\tilde{\pi}_{j+1}=j+1$. This means that $\tilde{\pi}_{j}\tilde{\pi}_{j+1}$ must be either the $34$ or the $12$ part of the $3412$ pattern. But if it were the $34$ part, then $\pi_{i}\pi_{j}x$ could be completed into two occurrences of $321$ in $\pi$ by setting $x$ to be the $1$ or the $2$ in the $3412$ occurrence in $\tilde{\pi}$ (these are the same in $\pi$ and $\tilde{\pi}$). Similarly, if $\tilde{\pi}_{j}\tilde{\pi}_{j+1}$ were the $12$ part of the $3412$ occurrence in $\tilde{\pi}$, then $y\pi_{j}\pi_{k}$ could be completed into two occurrences of $321$ in $\pi$ by setting $y$ to be the $3$ or the $4$ in the $3412$ occurrence in $\tilde{\pi}$. In both cases this contradicts the fact that $\pi$ contains a unique occurrence of $321$, and so the claim that $\tilde{\pi}$ avoids $3412$ is proved. We then define a permutation $\pi^{\prime}$, obtained from $\tilde{\pi}$ by changing the $4231$ pattern to a $3412$ pattern (on the same elements), and leaving all other elements unchanged. More formally, we have:
[…]
@@
The construction $\pi\mapsto\pi^{\prime}$ is illustrated in Figure 15 below. On the left (Figure 15(a)) we have the occurrence $\pi_{i},\pi_{j}=j,\pi_{k}$ of $321$ in the permutation $\pi\in B(n,2)$. On the right (Figure 15(b)) we have the occurrence $\pi^{\prime}_{i}=\pi_{j}+1=j+1,\pi^{\prime}_{j}=\pi_{i}+1,\pi^{\prime}_{j+1}=\pi_{k},\pi^{\prime}_{k+1}=\pi_{j}=j$ of $3412$ in $\pi^{\prime}$. The other dots of $\pi$ are unchanged in $\pi^{\prime}$, after the necessary shifting downwards and/or to the right to create a permutation.
[…]
@@
Let $\pi\in B(n,2)$. By construction $\pi^{\prime}$ is an $(n+1)$-permutation. We first show that $\pi^{\prime}\in B(n+1,3)$, using Proposition 4.12. We first refine the construction of Figure 15, using the mesh pattern of $\pi$ illustrated in Figure 12. For convenience, we represent this as a block decompositions of the permutations $\pi$ and $\pi^{\prime}$, as illustrated in Figure 16 below. The unlabelled blocks correspond to the shaded regions of the mesh patterns, and are therefore empty.